# Parsimonious Gaussian mixture models with piecewise-constant eigenvalue profiles

### v1

@@ 1 Introduction
Gaussian mixture models (GMMs) [49] are ubiquitous in statistical learning. Their probabilistic nature gives them many advantages against classical machine learning algorithms such as random forests [19], support vector machines [13], K-Means [44], HDBSCAN [47] and spectral clustering [51]. They notably enable uncertainty estimates, likelihood-based model selection, generative modeling and they can handle missing data. Gaussian-based models show good performances in low-to-moderate dimensions and are therefore heavily used in basic tasks such as classification or clustering. However, the overparameterization of their covariance matrices—$p(p+1)/2$ parameters in dimension $p$—is a clear issue in high dimensions, where data often lack, hence the need for simpler models.
@@
Many works have attempted to overcome the high-dimensional issues of GMMs (see for instance the Chapter 8 of [10] for a review). A widely-used heuristic is to reduce the dimension of the data—e.g. via principal component analysis [42]—before fitting the mixture model. This practical method may however obviously discard some important information for the subsequent task; it may for instance mix some clusters that were well separated in the original space [25]. Other methods—referred to as parsimonious—consider GMMs in the full data space but with fewer parameters. The early works of [18, 24, 6] consider GMMs with two simple parsimony-inducing principles: parameter sharing (e.g. constraining the mixture components to have the same covariance matrix) and covariance constraints (e.g. imposing diagonal or scalar covariance matrices). However, the former may not reduce enough the number of parameters in high dimensions, while the latter may reduce it too much. As a consequence, many intermediate parsimonious models later appeared, with low-rank [34, 50, 48], sparse [30, 22], Markov random field [32], ultrametric [26], or structured priors [37] for the covariance matrices, with an important impact [31, 57].
@@
Among the parsimonious GMM family, the celebrated mixtures of probabilistic principal component analyzers (MPPCA) [61] assume that each cluster lies around a low-dimensional affine subspace, following a multivariate Gaussian distribution with a specific eigenvalue profile—the smallest eigenvalues being constant and equal to the noise level. This model however overlooks an important difficulty which is the one of dimension selection; the learning of the mixture parameters indeed requires to prespecify the intrinsic dimensions of the clusters, which are often unknown. The high-dimensional data clustering (HDDC) [11] and high-dimensional discriminant analysis (HDDA) [12] models address this issue by automatically adjusting the clusters dimensions over the iterations of the optimization algorithm, based on Cattell’s scree test [21]. The theoretical guarantees of the learning algorithm are however not understood, and more practically, these models still require a hyperparameter as an input—the threshold for Cattell’s scree test—which has an important impact on the choice of intrinsic dimensions.
@@
Other methods for the automatic learning of intrinsic dimensions in parsimonious GMMs exist. A natural approach adopted in [48] is the grid search combined with statistical model selection criteria such as the Akaike information criterion (AIC) [3], the Bayesian information criterion (BIC) [56], or the integrated completed likelihood (ICL) [7]. While such a grid search is computationally expensive, few alternatives have been proposed: Ghahramani and Beal [33] used a sparsity-inducing prior together with variational inference. In contrast, the problem of estimating the intrinsic dimension of a single Gaussian has received a much wider attention, using for instance likelihood-based techniques [8], Bayesian model selection [14, 16], or approximate cross-validation [40].
@@
Meanwhile, the building blocks of low-rank GMMs—the PPCA models of [62]—have been extended into Gaussian models with piecewise-constant eigenvalue profiles under the name of principal subspace analysis (PSA) [59, 58]. The motivations are simple: if some sample eigenvalues are relatively close, then equalizing them slightly reduces the likelihood while drastically reducing the number of parameters. In that sense, the PSA models go beyond low-rank models like PPCA in terms of parsimony and flexibility—since any group of eigenvalues can be equalized and not just the smallest ones. The eigenvalues multiplicities are automatically learned by optimizing a penalized-likelihood criterion—e.g. the BIC. While PSA raises promising perspectives for parsimonious modeling and interpretability, it is certainly not suited for multimodal density modeling—which is the case of most real datasets—or classical learning tasks such as classification and clustering.
@@
In this paper, we address all the limits raised in the previous paragraphs at once with a new family of parsimonious GMMs termed mixtures of principal subspace analyzers (MPSA). These models consist in the multimodal extension of PSA models, or equivalently the parsimonious generalization of MPPCA models with arbitrary eigenvalues multiplicities. At fixed multiplicities, we propose an expectation–maximization (EM) algorithm [27] for the unsupervised learning of mixture parameters. When the multiplicities are not prespecified, we propose an efficient and provably-monotonous variant of the EM algorithm which automatically adjusts the groups of eigenvalues to equalize within the iterations. The underlying idea is that increasing the penalized-likelihood componentwise increases the mixture penalized-likelihood. Importantly, this algorithm is hyperparameter-free, in the sense that the groups of eigenvalues to equalize do not depend on any hyperparameter to tune—compared to low-rank models like MPPCA and HDDC. As we shall discuss it, we believe that our EM variant can apply far beyond the scope of the current paper, providing new principles for integrated parameter estimation and model selection. We finally perform several experiments (density estimation, clustering, single-image denoising) which all show the superior likelihood–parsimony tradeoffs achieved by our models with respect to the classical GMMs.
@@ 2 Principal subspace analysis
PSA is a Gaussian model with a constraint on the covariance eigenvalues multiplicities. It has been originally introduced to identify and overcome a “curse of isotropy” that arises in PCA when eigenvalues are relatively close. In this section, we summarize these ideas through the lens of parsimonious Gaussian modeling. More details can be found in the original paper [59], including illustrations of the generative models, their poset structure, and the impact on the PCA methodology.
@@
be the set of compositions of $p\in\mathbb{N}_{*}$ [5]. Let $\gamma\in\mathcal{C}(p)$, $q_{k}\coloneqq\sum_{k^{\prime}=1}^{k}\gamma_{k^{\prime}}$, and let us write $q\coloneqq q_{d-1}$. The PSA model of type $\gamma$ is defined as the parametric family of Gaussian densities ${}{p(\,x\mid\mu,\Sigma\,)=\mathcal{N}(\,x\mid\mu,\Sigma\,)}$, where $\mu\in\mathbb{R}^{p}$ is the mean and $\Sigma\in\mathrm{Sym}_{+}(p)$ is a covariance matrix with repeated eigenvalues of multiplicities $\gamma$. More precisely, one assumes that the covariance matrix takes the form ${}{\Sigma=\sum\nolimits_{k}\lambda_{k}\Pi_{k}}$, where ${}{\lambda_{k}\in\mathbb{R}_{+}}$ is the $k$-th largest eigenvalue (of multiplicity $\gamma_{k}$) and $\Pi_{k}\in\mathrm{Sym}(p)$ is the orthogonal projector (of rank $\gamma_{k}$) onto the associated eigenspace (of dimension $\gamma_{k}$). Such a set of covariance matrices with eigenvalue multiplicities $\gamma$ is written $\mathrm{Sym}_{+}(\gamma)$.
By definition, PSA models generalize the celebrated PPCA models of [62]—whose $q$ highest eigenvalues are (almost surely) distinct and $p-q$ remaining eigenvalues are equal to the noise level—with types of the form $\gamma=(1,\dots,1,p-q)$. [^1] We will sometimes use the notation $1^{a}$, as an abbreviation for $1,\dots,1$ (where 1 is repeated $a$ times). PSA models also generalize the isotropic PPCA models of [8]—whose eigenvalue profile forms two blocks, i.e. with types of the form $\gamma=(q,p-q)$.
PSA models are parameterized by the mean ${}{\mu\in\mathbb{R}^{p}}$ and the covariance $\Sigma\in\mathrm{Sym}_{+}(\gamma)$. One could therefore think that the number of parameters is ${}{p+{p(p+1)}/{2}}$. However, the repeated eigenvalues make the actual number of parameters much smaller. [59] show that the number of free parameters for a PSA model of type $\gamma$ is
@@
Given a dataset $\left(x_{i}\right)_{i=1}^{n}$ in $\mathbb{R}^{p}$, the PSA model of type $\gamma$ that is the most likely to have generated the data (i.e. the maximum likelihood estimate, MLE) can be computed in closed-form [59]. The optimal mean is the empirical mean $\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and the optimal covariance matrix $\hat{\Sigma}$ is a piecewise-constant version of the empirical covariance matrix, where the eigenvalues have been block-averaged according to $\gamma$. More precisely, if the empirical covariance $S\coloneqq\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}$ eigendecomposes as $S=\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$, where $\ell_{1}\geq\dots\geq\ell_{p}>0$ are the eigenvalues [^2] In the whole paper, for simplicity, we assume that the sample covariance matrices are always invertible. We make this assumption satisfied in practice by adding a covariance regularization term to the diagonal of the covariance matrices. and $v_{1}\perp\dots\perp v_{p}$ are the eigenvectors, then
@@
This model selection criterion results from an asymptotic approximation of the model evidence and it makes a tradeoff between model complexity and goodness-of-fit. Using the BIC, [59] get the important result that two adjacent sample eigenvalues $\ell_{j}$ and $\ell_{j+1}$ should be equalized when their relative eigengap $\delta(\ell_{j},\ell_{j+1})\coloneqq\frac{\ell_{j}-\ell_{j+1}}{\ell_{j}}$ is below a certain threshold $\delta(n)$ depending on the number of samples $n$ but not on the dimension $p$:
@@
Through extensive empirical experiments, [59] show that one rarely has enough samples to distinguish all the empirical eigenvalues, and that one should therefore transition from principal components to principal subspaces.
This important relative eigengap result suggests two simple methodologies to choose the groups of eigenvalues to equalize. The relative eigengap strategy consists in block-averaging the groups of eigenvalues whose relative eigengap is below the threshold $\delta(n)$. The hierarchical clustering strategy consists in a hierarchical clustering [66] of the eigenvalues, yielding a family of models $\mathcal{M}_{\mathrm{h}}(\ell)\in\mathcal{C}(p)^{p}$ of decreasing complexity, from which one chooses the model with the lowest BIC. The detailed algorithm is given in [59]. While the relative eigengap strategy is fast ($O(p)$) compared to the hierarchical clustering strategy ($O(p^{2})$), one can show that the former always selects a PSA model with greater or equal BIC, since the latter optimizes over a larger set. [^3] Indeed the (single-linkage) hierarchical clustering boils down to regrouping the eigenvalues sequentially according to their relative eigengap. Therefore, at some point, all the eigenvalue pairs below the relative eigengap threshold (8) are clustered, similarly as in the relative eigengap strategy. In other words, the family $\mathcal{M}_{\mathrm{h}}(\ell)\in\mathcal{C}(p)^{p}$ contains the model output by the relative eigengap strategy. The more recent eigengap sparsity of [58] relaxes the discrete BIC minimization problem into a continuous problem penalized by pairwise distances between sample eigenvalues [63].
@@ 3.1 Mixture density and parameters
Remark 1.
Similarly as PSA extends and unifies PPCA with other parsimonious Gaussian models (like IPPCA, spherical and full-Gaussian models), MPSA extends and unifies MPPCA with other classical parsimonious GMMs belonging to [11] and [48].
@@ 3.2 Inference and EM algorithm
$\hat{\theta}\in\Theta(\gamma)$triangleright mixture parameters
$y\leftarrow\operatorname{KMeans}(x,C)\in{[\![1,C]\!]}^{n}$triangleright K-Means init.
$t\leftarrow\left(\mathbf{1}_{y_{i}=c}\right)_{c,i}$triangleright posterior proba.
@@
     $\hat{\theta}\leftarrow\operatorname{M-step}(t,\gamma)$triangleright thm. 3
     $t\leftarrow\operatorname{E-step}(\hat{\theta})$triangleright thm. 2
@@ 3.3 EM complexity analysis
The K-Means clustering algorithm [44] involved in the EM initialization has time complexity $O(CnpT)$, with $T$ the number of iterations, which is by default at most $300$ in the scikit-learn implementation.
@@
A classical E-step would have time complexity $O(Cp^{3})$ due to the covariance inversion. However, the formulas that we derive in Theorem 2 have several computational advantages. First, they remove the necessity to numerically invert the covariance matrices—which can be pathological in high dimensions. Since the eigenvalues $\lambda_{ck}\in\mathbb{R}$ and the eigenvectors ${}{Q_{ck}\in\mathbb{R}^{p\times\gamma_{ck}}}$ are computed in the M-step, the computational complexity of the cost $K_{c}(x)$ is the one of ${}{\sum_{k}{\lambda_{ck}}^{-1}(x-\mu_{c})^{\top}\Pi_{ck}(x-\mu_{c})}$, which can be rewritten as ${}{\sum_{k}{\lambda_{ck}}^{-1}\|{Q_{ck}}^{\top}(x-\mu_{c})\|_{2}^{2}}$, that is $\sum_{k=1}^{d_{c}}O(\gamma_{ck}p)=O(p^{2})$. Therefore, the complexity of the E-step is $O(Cnp^{2})$. But this is not all. Since $(\Pi_{c1},\dots,\Pi_{cd_{c}})$ represents a flag of subspaces [69]—i.e. mutually-orthogonal subspaces spanning the whole space $\mathbb{R}^{p}$—then any orthogonal projection matrix $\Pi_{ck}$ can be expressed as $\Pi_{ck}=I-\sum_{k^{\prime}\neq k}\Pi_{ck^{\prime}}$. Notably, if one projector is of much greater rank than the other ones (for instance the last one $\Pi_{cd_{c}}$, accounting for the noise), then it can be replaced with lower dimensional terms in (13) and yield much more computationally-efficient formulas for the E-step. More precisely, if one has ${}{q_{c}\coloneqq p-\gamma_{cd_{c}}=O(1)}$, then the computational complexity of the cost function $K_{c}$ drops to $O(q_{c}p)=O(p)$. Eventually, the complexity of the E-step in high dimensions will more likely be $O(Cnpq)$, with ${}{q\coloneqq\max_{c\in[\![1,C]\!]}q_{c}}$.
@@
The M-step involves the eigenvalue decomposition of the local responsibility-weighted covariance matrices $S_{c}$. As currently implemented for simplicity, computing both the covariance matrices and their eigenvalue decompositions results in a complexity of $O(Cnp^{2})+O(Cp^{3})$. Therefore, even though the eigenvalue equalization of Theorem 3 is extremely simple, it does not remove the need for the costly eigenvalue decomposition. To make it more efficient in high dimensions, one could use classical numerical tricks such as the truncated singular value decomposition, which computes only a few leading eigenvectors. This would however require setting a bound on the intrinsic dimensions (i.e. $q_{c}\leq q_{\mathrm{max}}\,\forall c$), which can be restrictive depending on the context. We leave this more efficient implementation of the M-step to future research.
@@
In Algorithm 1, the mixture model types $\gamma$ are assumed to be prespecified. Unless they are known, a classical way to choose these hyperparameters is via cross-validation or statistical model selection techniques, implying running several EM algorithms [48]. The important issue here is that the number of candidate MPSA models is ${(2^{p-1})}^{C}$. Indeed, each independent mixture component is parameterized by its type $\gamma_{c}$ which can take $2^{p-1}$ values, since each adjacent-eigenvalue pair can be either equalized or not. Therefore, an exhaustive model selection would imply running $O(2^{Cp})$ EM algorithms, which becomes computationally prohibitive for $p\gtrapprox 10$—even though the EM itself may not be too much computationally costly after integrating the previously-exposed low-dimensional numerical tricks.
@@
Drawing from the HDDC models of [11], one could naturally be tempted to perform some kind of model selection at the component level (i.e. based on the PSA methodology of [59] presented in Section 2) at each EM iteration. This would add a negligible time complexity to Algorithm 1 ($\mathcal{O}(Cp^{2})$ for the hierarchical clustering strategy), instead of a multiplicative factor. However, such a variant of the EM algorithm would not enjoy any theoretical guarantee a priori. Consequently, we propose in the next section to reframe our intuitions as a penalized optimization problem, for which we show that a thoughtful variant of the EM algorithm yields a monotonous loss function.

### v2

@@ 1 Introduction
Gaussian mixture models (GMMs) [mclachlan_finite_2000] are ubiquitous in statistical learning. Their probabilistic nature gives them many advantages against classical machine learning algorithms such as random forests [breiman_random_2001], support vector machines [boser_training_1992], K-Means [lloyd_least_1982], HDBSCAN [mcinnes_hdbscan_2017] and spectral clustering [ng_spectral_2001]. They notably enable uncertainty estimates, likelihood-based model selection, generative modeling and they can handle missing data. Gaussian-based models show good performances in low-to-moderate dimensions and are therefore heavily used in basic tasks such as classification or clustering. However, the overparameterization of their covariance matrices—$p(p+1)/2$ parameters in dimension pp—is a clear issue in high dimensions, where data often lack. Hence, there is a need for simpler models.
@@
Many works have attempted to overcome the high-dimensional issues of GMMs (see for instance the Chapter 8 of [bouveyron_model-based_2019] for a review). A widely-used heuristic is to reduce the dimension of the data—e.g. via principal component analysis [jolliffe_principal_2002]—before fitting the mixture model. This practical method may however obviously discard some important information for the subsequent task; it may for instance mix some clusters that were well separated in the original space [chang_using_1983]. Other methods—referred to as parsimonious—consider GMMs in the full data space but with fewer parameters. The early works of [banfield_model-based_1993, celeux_gaussian_1995, bensmail_regularized_1996] consider GMMs with two simple parsimony-inducing principles: parameter sharing (e.g. constraining the mixture components to have the same covariance matrix) and covariance constraints (e.g. imposing diagonal or scalar covariance matrices). However, the former may not reduce enough the number of parameters in high dimensions, while the latter may reduce it too much. As a consequence, many intermediate parsimonious models later appeared, with low-rank [ghahramani_em_1996, mclachlan_modelling_2003, mcnicholas_parsimonious_2008], sparse [fop_model-based_2019, cappozzo_sparse_2025], Markov random field [finder_effective_2022], ultrametric [cavicchia_parsimonious_2024], or structured priors [heaps_structured_2024] for the covariance matrices, with an important impact [fraley_model-based_2002, scrucca_mclust_2016].
@@
Among the parsimonious GMM family, the celebrated mixtures of probabilistic principal component analyzers (MPPCA) [tipping_mixtures_1999] assume that each cluster lies around a low-dimensional affine subspace, following a multivariate Gaussian distribution with a specific eigenvalue profile—the smallest eigenvalues being constant and equal to the noise level. This model however overlooks an important difficulty which is the one of dimension selection; the learning of the mixture parameters indeed requires to prespecify the intrinsic dimensions of the clusters, which are often unknown. The high-dimensional data clustering (HDDC) [bouveyron_high-dimensional_2007] and high-dimensional discriminant analysis (HDDA) [bouveyron_high-dimensional_2007-1] models address this issue by automatically adjusting the clusters dimensions over the iterations of the optimization algorithm, based on Cattell’s scree test [cattell_scree_1966]. The theoretical guarantees of the learning algorithm are however not understood, and more practically, these models still require a hyperparameter as an input—the threshold for Cattell’s scree test—which has an important impact on the choice of intrinsic dimensions.
@@
Other methods for the automatic learning of intrinsic dimensions in parsimonious GMMs exist. A natural approach adopted in [mcnicholas_parsimonious_2008] is the grid search combined with statistical model selection criteria such as the Akaike information criterion (AIC) [akaike_new_1974], the Bayesian information criterion (BIC) [schwarz_estimating_1978], or the integrated completed likelihood (ICL) [biernacki_assessing_2000]. While such a grid search is computationally expensive, few alternatives have been proposed: ghahramani_variational_1999 used a sparsity-inducing prior together with variational inference. In contrast, the problem of estimating the intrinsic dimension of a single Gaussian has received a much wider attention, using for instance likelihood-based techniques [bouveyron_intrinsic_2011], Bayesian model selection [bishop_bayesian_1998, bouveyron_exact_2020], or approximate cross-validation [josse_selecting_2012].
@@
Meanwhile, the building blocks of low-rank GMMs—the PPCA models of [tipping_probabilistic_1999]—have been extended into Gaussian models with piecewise-constant eigenvalue profiles under the name of principal subspace analysis (PSA) [szwagier_curse_2025, szwagier_eigengap_2025]. The motivations are simple: if some sample eigenvalues are relatively close, then equalizing them slightly reduces the likelihood while drastically reducing the number of parameters. In that sense, the PSA models go beyond low-rank models like PPCA in terms of parsimony and flexibility—since any group of eigenvalues can be equalized and not just the smallest ones. The eigenvalues multiplicities are automatically learned by optimizing a penalized log-likelihood criterion—e.g. the BIC. While PSA raises promising perspectives for parsimonious modeling and interpretability, it is certainly not suited for multimodal density modeling—which is the case of most real datasets—or classical learning tasks such as classification and clustering.
@@
In this paper, we address all the limits raised in the previous paragraphs at once with a new family of parsimonious GMMs termed mixtures of principal subspace analyzers (MPSA). These models consist in the multimodal extension of PSA models, or equivalently the parsimonious generalization of MPPCA models with arbitrary eigenvalues multiplicities. At fixed multiplicities, we propose an expectation–maximization (EM) algorithm [dempster_maximum_1977] for the unsupervised learning of mixture parameters. When the multiplicities are not prespecified, we propose an efficient and provably-monotonous variant of the EM algorithm which automatically adjusts the groups of eigenvalues to equalize within the iterations. The underlying idea is that increasing the penalized log-likelihood componentwise increases the mixture penalized log-likelihood. Importantly, this algorithm is hyperparameter-free, in the sense that the groups of eigenvalues to equalize do not depend on any hyperparameter to tune—compared to low-rank models like MPPCA and HDDC. As we shall discuss it, we believe that our EM variant can apply far beyond the scope of the current paper, providing new principles for integrated parameter estimation and model selection. We finally perform several experiments (density estimation, clustering, single-image denoising) which all show the superior likelihood–parsimony tradeoffs achieved by our models with respect to the classical GMMs.
@@ 2 Principal subspace analysis
PSA is a Gaussian model with a constraint on the covariance eigenvalues multiplicities. It has been originally introduced to identify and overcome a “curse of isotropy” that arises in PCA when eigenvalues are relatively close. In this section, we summarize these ideas through the lens of parsimonious Gaussian modeling. More details can be found in the original paper [szwagier_curse_2025], including illustrations of the generative models, their poset structure, and the impact on the PCA methodology.
@@
be the set of compositions of $p\in\mathbb{N}_{*}$ [bergeron_standard_1995]. Let $\gamma\in\mathcal{C}(p)$, $q_{k}\coloneqq\sum_{k^{\prime}=1}^{k}\gamma_{k^{\prime}}$, and let us write $q\coloneqq q_{d-1}$. The PSA model of type $\gamma$ is defined as the parametric family of Gaussian densities $p(\,x\mid\mu,\Sigma\,)=\mathcal{N}(\,x\mid\mu,\Sigma\,)$}, where $\mu\in\mathbb{R}^{p}$ is the mean and $\Sigma\in\mathrm{Sym}_{+}(p)$ is a covariance matrix with repeated eigenvalues of multiplicities $\gamma$. More precisely, one assumes that the covariance matrix takes the form $\Sigma=\sum\nolimits_{k}\lambda_{k}\Pi_{k}$}, where $\lambda_{k}\in\mathbb{R}_{+}$} is the $k$-th largest eigenvalue (of multiplicity $\gamma_{k}$) and $\Pi_{k}\in\mathrm{Sym}(p)$ is the orthogonal projector (of rank $\gamma_{k}$) onto the associated eigenspace (of dimension $\gamma_{k}$). Such a set of covariance matrices with eigenvalue multiplicities $\gamma$ is written $\mathrm{Sym}_{+}(\gamma)$.
By definition, PSA models generalize the celebrated PPCA models of [tipping_probabilistic_1999]—whose $q$ highest eigenvalues are (almost surely) distinct and pqp-q remaining eigenvalues are equal to the noise level—with types of the form $\gamma=(1,\dots,1,p-q)$. [^1] We will sometimes use the notation $1^{a}$, as an abbreviation for $1,\dots,1$ (where 1 is repeated $a$ times). PSA models also generalize the isotropic PPCA models of [bouveyron_intrinsic_2011]—whose eigenvalue profile forms two blocks, i.e. with types of the form $\gamma=(q,p-q)$.
PSA models are parameterized by the mean $\mu\in\mathbb{R}^{p}$} and the covariance $\Sigma\in\mathrm{Sym}_{+}(\gamma)$. One could therefore think that the number of parameters is p+p(p+1)/2~{p+{p(p+1)}/{2}}. However, the repeated eigenvalues make the actual number of parameters much smaller. [szwagier_curse_2025] show that the number of free parameters for a PSA model of type $\gamma$ is
@@
Given a dataset $\left(x_{i}\right)_{i=1}^{n}$ in $\mathbb{R}^{p}$, the PSA model of type $\gamma$ that is the most likely to have generated the data (i.e. the maximum likelihood estimate, MLE) can be computed in closed-form [szwagier_curse_2025]. The optimal mean is the empirical mean $\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and the optimal covariance matrix $\hat{\Sigma}$ is a piecewise-constant version of the empirical covariance matrix, where the eigenvalues have been block-averaged according to $\gamma$. More precisely, if the empirical covariance $S\coloneqq\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}$ eigendecomposes as $S=\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$, where $\ell_{1}\geq\dots\geq\ell_{p}>0$ are the eigenvalues [^2] In the whole paper, for simplicity, we assume that the sample covariance matrices are always invertible. We make this assumption satisfied in practice by adding a covariance regularization term to the diagonal of the covariance matrices. and $v_{1}\perp\dots\perp v_{p}$ are the eigenvectors, then
@@
This model selection criterion results from an asymptotic approximation of the model evidence and it makes a tradeoff between model complexity and goodness-of-fit. Using the BIC, [szwagier_curse_2025] get the important result that two adjacent sample eigenvalues $\ell_{j}$ and $\ell_{j+1}$ should be equalized when their relative eigengap $\delta(\ell_{j},\ell_{j+1})\coloneqq\frac{\ell_{j}-\ell_{j+1}}{\ell_{j}}$ is below a certain threshold $\delta(n)$ depending on the number of samples $n$ but not on the dimension $p$:
@@
Through extensive empirical experiments, [szwagier_curse_2025] show that one rarely has enough samples to distinguish all the empirical eigenvalues, and that one should therefore transition from principal components to principal subspaces.
This important relative eigengap result suggests two simple methodologies to choose the groups of eigenvalues to equalize. The relative eigengap strategy consists in block-averaging the groups of eigenvalues whose relative eigengap is below the threshold $\delta(n)$. The hierarchical clustering strategy consists in a hierarchical clustering [ward_hierarchical_1963] of the eigenvalues, yielding a family of models $\mathcal{M}_{\mathrm{h}}(\ell)\in\mathcal{C}(p)^{p}$ of decreasing complexity, from which one chooses the model with the lowest BIC. The detailed algorithm is given in [szwagier_curse_2025]. While the relative eigengap strategy is fast ($O(p)$) compared to the hierarchical clustering strategy ($O(p^{2})$), one can show that the former always selects a PSA model with greater or equal BIC, since the latter optimizes over a larger set. [^3] Indeed the (single-linkage) hierarchical clustering boils down to regrouping the eigenvalues sequentially according to their relative eigengap. Therefore, at some point, all the eigenvalue pairs below the relative eigengap threshold (8) are clustered, similarly to the relative eigengap strategy. In other words, the family $\mathcal{M}_{\mathrm{h}}(\ell)\in\mathcal{C}(p)^{p}$ contains the model output by the relative eigengap strategy. The more recent eigengap sparsity of [szwagier_eigengap_2025] relaxes the discrete BIC minimization problem into a continuous problem penalized by pairwise distances between sample eigenvalues [tyler_lassoing_2020].
@@ 3.1 Mixture density and parameters
Remark 1 (Generalization of parsimonious GMMs).
As PSA extends and unifies PPCA with other parsimonious Gaussian models (like IPPCA, spherical and full-Gaussian models), MPSA extends and unifies MPPCA with other classical parsimonious GMMs belonging to [bouveyron_high-dimensional_2007] and [mcnicholas_parsimonious_2008].
@@ 3.2 Inference and EM algorithm
$\hat{\theta}\in\Theta(\gamma)$ $\triangleright$ mixture parameters
$y\leftarrow\operatorname{KMeans}(x,C)\in{[\![1,C]\!]}^{n}$ $\triangleright$ K-Means init.
$t\leftarrow\left(\mathbf{1}_{y_{i}=c}\right)_{c,i}$ $\triangleright$ posterior proba.
@@
  $\hat{\theta}\leftarrow\operatorname{M-step}(t,\gamma)$ $\triangleright$ thm. 3
[…]
@@ 3.3 EM complexity analysis
The K-Means clustering algorithm [lloyd_least_1982] involved in the EM initialization has time complexity $O(CnpT)$, with $T$ the number of iterations, which is by default at most $300$ in the scikit-learn implementation [pedregosa_scikit-learn_2011].
@@
A classical E-step would have time complexity $O(Cp^{3})$ due to the covariance inversion. However, the formulas that we derive in Theorem 2 have several computational advantages. First, they remove the necessity to numerically invert the covariance matrices—which can be pathological in high dimensions. Since the eigenvalues $\lambda_{ck}\in\mathbb{R}$ and the eigenvectors $Q_{ck}\in\mathbb{R}^{p\times\gamma_{ck}}$} are computed in the M-step, the computational complexity of the cost $K_{c}(x)$ is the one of $\sum_{k}{\lambda_{ck}}^{-1}(x-\mu_{c})^{\top}\Pi_{ck}(x-\mu_{c})$}, which can be rewritten as $\sum_{k}{\lambda_{ck}}^{-1}\|{Q_{ck}}^{\top}(x-\mu_{c})\|_{2}^{2}$}, that is $\sum_{k=1}^{d_{c}}O(\gamma_{ck}p)=O(p^{2})$. Therefore, the complexity of the E-step is $O(Cnp^{2})$. But this is not all. Since $(\Pi_{c1},\dots,\Pi_{cd_{c}})$ represents a flag of subspaces [ye_optimization_2022]—i.e. mutually-orthogonal subspaces spanning the whole space $\mathbb{R}^{p}$—then any orthogonal projection matrix $\Pi_{ck}$ can be expressed as $\Pi_{ck}=I-\sum_{k^{\prime}\neq k}\Pi_{ck^{\prime}}$. Notably, if one projector is of much greater rank than the other ones (for instance the last one $\Pi_{cd_{c}}$, accounting for the noise), then it can be replaced with lower dimensional terms in (13) and yield much more computationally-efficient formulas for the E-step. More precisely, if one has $q_{c}\coloneqq p-\gamma_{cd_{c}}=O(1)$}, then the computational complexity of the cost function $K_{c}$ drops to $O(q_{c}p)=O(p)$. Eventually, the complexity of the E-step in high dimensions will more likely be $O(Cnpq)$, with $q\coloneqq\max_{c\in[\![1,C]\!]}q_{c}$}.
@@
The M-step involves the eigenvalue decomposition (EVD) of the local responsibility-weighted covariance matrices $S_{c}$. As currently implemented for simplicity, computing both the covariance matrices and their EVDs results in a complexity of $O(Cnp^{2})+O(Cp^{3})$. Therefore, even though the eigenvalue equalization of Theorem 3 is extremely simple, it does not remove the need for the costly EVD. To make it more efficient in high dimensions, one can use classical numerical tricks such as the truncated singular value decomposition (SVD), which computes only a few leading eigenvectors. This perspective is investigated in Section B.4.
@@
In Algorithm 1, the mixture model types $\gamma$ are assumed to be prespecified. Unless they are known, a classical way to choose these hyperparameters is via cross-validation or statistical model selection techniques, implying running several EM algorithms [mcnicholas_parsimonious_2008]. The important issue here is that the number of candidate MPSA models is ${(2^{p-1})}^{C}$. Indeed, each independent mixture component is parameterized by its type $\gamma_{c}$ which can take $2^{p-1}$ values, since each adjacent-eigenvalue pair can be either equalized or not. Therefore, an exhaustive model selection would imply running $O(2^{Cp})$ EM algorithms, which becomes computationally prohibitive for $p\gtrapprox 10$—even though the EM itself may not be too much computationally costly after integrating the previously-exposed low-dimensional numerical tricks.
@@
Drawing from the HDDC models of [bouveyron_high-dimensional_2007], one could naturally be tempted to perform some kind of model selection at the component level (i.e. based on the PSA methodology of [szwagier_curse_2025] presented in Section 2) at each EM iteration. This would add a negligible time complexity to Algorithm 1 ($\mathcal{O}(Cp^{2})$ for the hierarchical clustering strategy), instead of a multiplicative factor. However, such a variant of the EM algorithm would not enjoy any theoretical guarantee a priori. Consequently, we propose in the next section to reframe our intuitions as a penalized optimization problem, for which we show that a thoughtful variant of the EM algorithm yields a monotonous loss function.